OpenStudy (anonymous):

Find the volume of the solid that lies within both the cylinder x^2 + y^2 = 4 and the sphere x^2 + y^2 + z^2 = 16. please help!!!! will give medal and fan

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

\[ 8 \times \int_0^2\int_0^{\sqrt{4 - y^2 }} \sqrt{16 - x^2 - y^2} dy dx\]

OpenStudy (experimentx):



OpenStudy (experimentx):